# Design Principles of Zero-Shot Self-Supervised Unknown Emitter Detectors

### v1

@@ VI-D The impact of the number of clusters in DMM
Figure 8 shows the PCA performance in terms of ROC-AUC, NMI, and F1 metrics for varying numbers of clusters in the decision-making step on the ORACLE and WiSig datasets. The results reveal that ROC-AUC and F1 reach a plateau, with only minor changes once the number of clusters becomes sufficiently large, while NMI consistently decreases. The Decreasing of the NMI metric does not mean the decreasing of clusters assignment [mahmoudi2024proof], because NMI also depends on the number of train emitters. These results indicate that the number of clusters required to reach stability is dataset-dependent. For the WiSig dataset with 6 emitters, stability is reached at approximately 80 clusters, whereas for the ORACLE dataset with 16 emitters, the stable region begins from approximately 200 clusters.
Figure 8: Performance of PCA with feature size 20 on WiSig (blue) and ORACLE (orange) datasets over different numbers of clusters. 2D-Constellation transformation has been applied to samples from the ORACLE dataset.
@@
Figure 9: Interpretability analysis for WiSig dataset
@@ VI-E Interpretability evaluation
At last, we evaluate the implementation approaches to the proposed concept in terms of interpretability. CNNs incorporate several techniques for interpreting model decisions. One widely used method is LIME (Local Interpretable Model-agnostic Explanations), which approximates the model’s behavior around a specific input sample using a linear model:
@@
Here, $x$ is the input sample, and the matrix $\mathbf{W}$ captures the local linear approximation of the model. LIME can further be used to analyze how specific inputs influence the sample’s position in the embedding space, or serve as a locally fully explainable method—especially after pruning non-essential weights from $\mathbf{W}$.
For the WiSig dataset, this concept is illustrated in Fig.9(a), where a linear layer approximates the CNN’s behavior in the vicinity of a data point. The corresponding distribution of weights from these locally fitted linear models is shown in Fig.9(b). However, due to the inherent non-linearity of CNNs, such local interpretability methods cannot generalize across the entire input space.
In contrast, Kolmogorov–Arnold Networks provide global interpretability. As depicted in Fig.9(c), each component of the 2D embedding space is connected to the inputs via known spline functions. The most important connections (black lines) can be approximated using symbolic regression, while less important ones are shown in gray. The possible symbolic regression of spline functions is shown in Fig.9(d).
This structure enables tracing how input perturbations affect internal representations. The distribution of input importance across KANs is illustrated in Fig.9(e), providing a more holistic view of feature influence.
@@ VII Conclusions
This paper presents a comprehensive analysis of the design space for Unknown Emitter Detection (UED) systems. We explore the design dimensions across various transmission scenarios (same and different messages), 1D raw I/Q and novel 2D-Constellation data modalities, feature learning modules (CNN and KAN), and machine learning paradigms (DC, AE, CL). Our theoretical and experimental findings reveal the following key insights:
  • The 2D-Constellation data modality for zero-shot self-supervised UED achieves up to a 40% performance improvement compared to the conventional raw I/Q input format under the different-messages scenario. To the best of our knowledge, this is the first work to address the challenge of differing messages in SEI/UED tasks.
  • The KAN-based interpretable feature learning module for zero-shot self-supervised UED delivers comparable or superior performance relative to conventional CNN architectures.
  • A SVD-based initialization strategy for feature learning modules enhances performance by up to 40% compared to random initialization when applied to the sparse 2D-Constellation input format.
In summary, the results for zero-shot self-supervised UED demonstrate that incorporating KAN-based feature extractors and interpretable representations at both the decision-making and embedding levels enables the proposed model to generalize effectively to unseen emitters and variable signal conditions, a capability not previously demonstrated in the literature.
VIII Acknoweledgements
This work was supported in part by the European Union’s Horizon Europe research and innovation programme under the NANCY project (No. 101096456) and in part by the Slovenian Research and Innovation Agency under the grant P2-0016.

### v2

@@ VI-D The impact of the number of clusters in DMM
Figure 8 shows the DMM’s sensitivity to the number of selected clusters using the PCA baseline and measured with ROC-AUC, NMI, and F1 metrics on the ORACLE and WiSig datasets. The results reveal that ROC-AUC and F1 reach a plateau, with only minor changes once the number of clusters becomes sufficiently large, while NMI consistently decreases. The decrease in the NMI metric does not mean the decrease in cluster assignment [mahmoudi2024proof], because NMI also depends on the number of train emitters. These results indicate that the number of clusters required to reach stability is dataset-dependent. For the WiSig dataset with six emitters, stability is reached at approximately 80 clusters, whereas for the ORACLE dataset with 16 emitters, the stable region begins at approximately 200 clusters.
@@
Figure 9: Model interpretability insights for WiSig dataset
@@ VI-E Interpretability evaluation
Ultimately, we assess the implementation approaches to the proposed concept in terms of interpretability. CNNs incorporate several techniques for interpreting model decisions. One widely used method is LIME, which approximates the model’s behavior around a specific input sample using a linear model:
@@
Here, $x$ is the input sample, and the matrix $\mathbf{W}$ captures the local linear approximation of the model. LIME can further be used to analyze how specific inputs influence the sample’s position in the embedding space, or serve as a locally fully explainable method, especially after pruning non-essential weights from $\mathbf{W}$.
For the WiSig dataset, this concept is illustrated in Figure 9(a), where a linear layer approximates the CNN’s behavior in the vicinity of a data point. The corresponding distribution of weights from these locally fitted linear models is shown in Figure 9(b). However, due to the inherent non-linearity of CNNs, such local interpretability methods cannot generalize across the entire input space. In contrast, Kolmogorov–Arnold Networks provide global interpretability. As depicted in Figure 9(c), each component of the 2D embedding space is connected to the inputs via known spline functions. The most important connections (black lines) can be approximated using symbolic regression, while less important ones are shown in gray. The possible symbolic regression of spline functions is shown in Figure 9(d). This structure enables tracing how input perturbations affect internal representations. The distribution of input importance across KAN s is illustrated in Figure 9(e), providing a more holistic view of feature influence.
@@ VII Conclusions
This paper presents a comprehensive analysis of the design space for Unknown Emitter Detection. We explore the key design dimensions with respect to SMS and DMS. Through the lens of the ML workflow, we investigate the impact of (a) data modality, (b) feature learning module, (c) machine learning approach, and (d) decision-making module. Our theoretical and experimental results reveal several key insights: 1) Under realistic DMS conditions, using a 2D constellation data modality improves performance by up to 20 p.p. in ROC-AUC compared to raw I/Q data. 2) KAN provides an interpretable architecture and performs comparably to CNNs, a “black-box” alternative, across both SMS and DMS settings. 3) For DMS, the best-performing configuration combines an SVD-initialized KAN with a deep clustering approach and 2D constellation data. This setup outperforms a standard KAN (with all other workflow components held constant) by up to 20 p.p. in ROC-AUC. 4) In analyzing the decision-making module, we identify the optimal number of clusters required for environments with varying numbers of known emitters. In summary, the results for UED demonstrate that incorporating KAN-based feature extractors and interpretable representations at both the Decision Making Module and Feature Learning Module enables the proposed model to generalize effectively to unseen emitters and varying signal conditions, a capability not previously shown in the literature.